Sally travels 15m south 20m west find the displacement and distance?
1
Expert's answer
2020-03-13T11:04:24-0400

The distance


"d=15+20=35\:\rm m."

The displacement


"s=\sqrt{15^2+20^2}=25\:\rm m."